# A comparison of pre-existing $\Lambda$CDM predictions with the abundance of JWST galaxies at high redshift

### v1

@@ 2.1.3 Dust model
Dust attenuates the luminosities of massive galaxies and is an important consideration when predicting the luminosity function of galaxies observed in different wavebands. galform includes a self-consistent dust model that calculates the absorption of radiation by dust in UV, optical, and near-IR bands, as well as the re-radiation of this energy at IR and sub-mm wavelengths. The model assumes a two-phase dust medium, with dust in molecular clouds embedded in a diffuse dust component. The absorption and re-radiation of energy by the two dust components are calculated separately and, thus, the temperature of the different dust components can vary. Some simplifying assumptions are made: (i) a single dust temperature is assumed for each dust component (cloud or diffuse); (ii) temperature fluctuations for small grains are ignored; (iii) the dust opacity is assumed to be a power law at IR/sub-mm wavelengths, and (iv) polycyclic aromatic hydrocarbon (PAH) features are ignored. We refer readers to Lacey et al. (2016) for a more detailed description of the dust absorption and emission model.
@@ 2.2 Model parameters
Following the philosophy behind galform, as set out in Cole et al. (2000), Lacey et al. (2016), and other galform papers, we calibrate the model parameters so as to reproduce a small set of observational constraints, mainly at low redshift, including: (i) the $b_{J}-$ and $K-$band luminosity functions at $z=0$; (ii) the morphological fractions as a function of $r-$band luminosity of galaxies at $z=0$; (iii) the black hole–bulge mass relation at $z=0$; (iv) the number counts and redshift distribution of sub-millimetre galaxies; and (v) UV galaxy luminosity functions at $z=3$ and $z=6$.
Here, we calculate the same models as Cowley et al. (2018) (based on Lacey et al. 2016). We adopt the same two SN feedback models, STD and EvolFB, described in Section 2.1.1, retaining the same values of all the model parameters. For example, we did not change the parameters of the EvolFB model to reflect a recent change in the inferred redshift of reionization from Planck data, from $z_{\rm re}\gtrsim 8.5$ (Planck Collaboration et al. 2016) to $z_{\rm re}\sim 7-8$ (Planck Collaboration et al. 2020).
The main properties of the local galaxy population and of the submillimetre galaxies used to calibrate the model parameters for both the STD and EvolFB models applied to the P–Millennium simulation are shown in Fig. 2. All other properties of the resulting galaxy population are predictions of the model. In particular, the UV luminosity function of galaxies at the redshifts relevant to this work, $z=7-16$, was not used to calibrate the model. As we can see from Fig. 2, both the STD model and the EvolFB model show remarkable agreement with the observations at lower redshifts, which is consistent with the results of Lacey et al. (2016) and Cowley et al. (2019) for the STD model and of Hou et al. (2016) for the EvolFB model. The STD model is seen to be quite close to the EvolFB model at $z=0$ as the evolving SN feedback converges to the standard case below $z=4$ (Hou et al., 2016).
Figure 3: galform predictions for – upper left: the stellar mass function; upper right: the star-formation rate (SFR) distribution function; lower left: the correlation between stellar mass and halo mass, with error bars indicating the 16th to 84th percentile (±$\pm 1\sigma$) range; lower right: the correlation between UV magnitude and halo mass with error bars indicating the 16th to 84th percentile (±$\pm 1\sigma$) range. In the bottom two panels, only central galaxies with stellar mass $\log\,M_{\ast}/h^{-1}\mathrm{M_{\odot}}>6.5$ and UV magnitude $M_{\rm UV}<-15$ are included. In each panel, the results from $z=7$ to $z=16$ are shown using different colours, as indicated by the key. The STD model and EvolFB model, respectively, are shown with dashed and solid curves. Both models exhibit a similar level of scatter; for clarity, we only show the error bars for the EvolFB model in the bottom two panels.
@@
Before comparing galform predictions with the observational data from JWST, in Fig. 3, we present the predictions of the galform model for (i) the stellar mass function (SMF), (ii) the star-formation rate (SFR) distribution function, (iii) the correlation between halo mass and stellar mass and (iv) the correlation between halo mass and UV magnitude for both the STD and EvolFB models, from $z=7$ to $z=16$. We define:
@@
We can see from the upper left panel of Fig. 3 that the stellar mass function evolves significantly from $z=7$ to $z=16$. The SMF at the lower redshifts shows both a higher number density at fixed stellar mass and extends to larger masses. Across the whole redshift and stellar mass ranges, the EvolFB model predicts a higher galaxy abundance than the STD model. This is a natural consequence of the redshift-dependent supernova feedback model, in which the feedback becomes weaker at higher redshifts ($z>4$), allowing star formation to continue more efficiently.
The SFR distribution function (upper right panel) also shows significant evolution, with the SFR increasing, on average, towards lower redshifts. The SFR distribution function in the EvolFB model has a similar shape to that in the STD model at $z\lesssim 10$, but at $z\gtrsim 12$, the distribution in the EvolFB model peaks at a slightly higher SFR. In the STD model, the SFR peaks at around $0.1\,h^{-1}\,\mathrm{M_{\odot}\,yr^{-1}}$ at all redshifts, while it peaks at around $1\,h^{-1}\,\mathrm{M_{\odot}\,yr^{-1}}$ in the EvolFB model at $z\gtrsim 12$. The highest SFRs at $z\gtrsim 12$ hardly exceed $30\,h^{-1}\,\mathrm{M_{\odot}\,yr^{-1}}$ in either model. These observations are consistent with the results of Cowley et al. (2018, fig. 3), which provides a reassuring test of reproducibility in the galform model.
@@
Interestingly, both the STD and EvolFB models show a nearly power-law relation between halo mass and stellar mass, without significant redshift evolution. At fixed halo mass, the EvolFB model predicts more massive galaxies than the STD model, which is consistent with the higher abundance of galaxies at a given stellar mass in the EvolFB model seen in the upper left panel of Fig. 3. Unlike the $M_{\rm halo}-M_{\ast}$ relation, the $M_{\rm halo}-M_{\rm UV}$ relation exhibits a strong redshift dependence in both models: low redshift galaxies have brighter $M_{\rm UV}$ with increasing halo mass, while high redshift ($z\sim 14-16$) galaxies have a flatter $M_{\rm halo}-M_{\rm UV}$ relation. At fixed halo mass, high redshift galaxies appear to be brighter in the UV band than low redshift galaxies. This implies that the “bright” galaxies observed at high redshifts are not “massive” galaxies, contrary to what is often stated in the observational literature (e.g. Harikane et al., 2023; Labbé et al., 2023).
@@
Fig. 4 presents the dust-extincted UVLFs (Section 2.1.3) predicted at $z=7-16$ by galform based on the STD model (dashed curves) and the EvolFB model (solid curves). These are compared with observational data at the corresponding redshifts, including the new JWST observations, as well as earlier estimates using other telescopes such as the Hubble Space Telescope (HST). We use different symbols to denote different observational datasets: (i) UVLFs from spectroscopically confirmed ( $z_{\rm spec}$) high-$z$ candidates from JWST are shown using filled green triangles (Harikane et al., 2024), (ii) UVLFs derived using photometric redshifts $(z_{\rm photo})$ from JWST observations are shown using open green triangles (Adams et al., 2023; Bouwens et al., 2023; Castellano et al., 2023; Donnan et al., 2023; Finkelstein et al., 2023; Harikane et al., 2023; Leung et al., 2023; Morishita & Stiavelli, 2023; Pérez-González et al., 2023; Robertson et al., 2023; Willott et al., 2023), while (iii) UVLFs constructed from earlier, pre-JWST $z_{\rm photo}$ data are indicated by open black triangles (Bouwens et al., 2011, 2015, 2021; Bowler et al., 2014, 2020; Finkelstein et al., 2015; McLure et al., 2013; Morishita et al., 2018; Oesch et al., 2014, 2018; Schenker et al., 2013; Stefanon et al., 2019).
@@
As we can see from Fig. 4, the EvolFB model always predicts a higher abundance of galaxies at the faint end across the whole redshift range ($z=7-16$). This is as expected because the evolving supernova feedback model gives weaker feedback at low masses as well as at higher redshifts (Hou et al., 2016). Below $z\sim 12$, the EvolFB model predicts slightly fewer bright galaxies than the STD model. We note that the discrepancy between the STD and EvolFB models at the bright end at $z\lesssim 12$ is slightly larger than in Cowley et al. (2018). This is because of the difference in the dust attenuation models used in this work and in Cowley et al. (2018). Our main conclusions are unaffected by this difference. We note that the effects of dust extinction on the bright end of the UVLF are predicted to be large even at the high redshifts shown here, but we return to this issue in Section 4.1.
@@
At higher redshifts, however, a discrepancy between the models (with dust extinction) and the observational data becomes apparent. At $z=12$, the STD model underpredicts the abundance of galaxies over the entire $M_{\rm UV}$ range, while the EvolFB model agrees with the data at the faint end ($M_{\rm UV}\gtrsim-18.5$) but underpredicts the galaxy abundance at the bright end ($M_{\rm UV}\lesssim-19$) by $1-2$ dex. At $z=14$, the predictions of the EvolFB model are within the measurement uncertainty of the data at $M_{\rm UV}>-18$, but are significantly lower at the brighter end ($M_{\rm UV}<-18$) by $2-3$ dex. These results agree with the findings of Kannan et al. (2023) who compared UVLFs from JWST with a suite of simulations based on the IllustrisTNG galaxy formation model. Our results indicate that a revision of the original galform model is required to match the observational data. This is the focus of the next section.
@@ 4 Updating galform
As we have just seen, the predictions formulated by Cowley et al. (2018), based on two different variations of the galform semi-analytic model within the $\Lambda$CDM framework (the STD and EvolFB models; Section 2), are in good agreement with the latest JWST UVLF data for $z\lesssim 10$. However, the models underpredict the abundance of galaxies at the bright end of the UVLF ($M_{\rm UV}<-18$) at higher redshifts ($z\sim 12-14$) when dust extinction is included. This discrepancy could be due to several factors. In this section, we explore the sensitivity of the predicted UVLFs to the details of specific physical components of the model – most notably, the dust model and the IMF (described in Section 2). We investigate whether reasonable adjustments of the model parameters make it possible to bring the galform predictions into closer agreement with the observed UVLFs at $z\gtrsim 12$.
@@ 4.1 The effect of dust attenuation
A possible way to resolve the difference between our model predictions and the observed UVLFs at high redshifts is to alter the importance of dust attenuation, since dust in galaxies absorbs light and thus reduces the number counts of bright galaxies. As a limiting case, we begin by turning off the dust attenuation in galform altogether (for both the STD and EvolFB models) and calculate dust-free UVLFs at $z=7-16$; these predictions are also shown in Fig. 4 (blue curves).
@@
There is some observational evidence that dust may be less important at high redshift than in the local Universe. Several studies (e.g. Vladilo et al., 2011; De Cia et al., 2013, 2016; Wiseman et al., 2017) have reported a metallicity-dependent dust-to-metal ratio (DTM), which is about 50% lower at metallicity $Z=0.1Z_{\odot}$ than at solar metallicity, $Z_{\odot}$ (see Maiolino & Mannucci 2019 for a review). On the other hand, Wiseman et al. (2017) reported that they did not find any obvious redshift-dependence of the DTM. However, this conclusion is based on small galaxy samples up to a redshift of only $z\sim 5$. In galform, a constant DTM is assumed with a value of 0.334 (Lacey et al., 2016) in order to match the dust-to-gas ratio in the solar neighbourhood, $6.7\times 10^{-3}$ for $Z_{\odot}=0.02$ (Silva et al., 1998). If the reports of a lower DTM ratio at lower metallicity are correct, galform would overestimate the dust content at low metallicity and, consequently, that of galaxies at high redshifts.
Another, more radical, possibility is that dust may not have had time to form, at least, in some of the high redshift galaxies. The details of dust formation and destruction mechanisms are very uncertain even at low redshift but, from analytical and observational considerations, Mattsson & Andersen (2012) concluded that dust growth in the interstellar medium is the most important dust formation mechanism. They proposed a dust grain growth timescale, $\tau_{\rm growth}$, which is found to increase with decreasing metallicity and can be as long as $\gtrsim 1\,\mathrm{Gyr}$ for the lowest metallicity galaxies (Galliano et al., 2021). Since redshift $z=14$ corresponds to only $\sim 300\,\mathrm{Myrs}$ after the Big Bang, galaxies at this redshift may not have had enough time to make a significant amount of dust. Schneider & Maiolino (2024) proposed a parametrization of $\tau_{\rm growth}$ which depends on the distribution of grain sizes, and on the density, temperature and metallicity of the interstellar medium. This model also shows a negative correlation with metallicity, similar to that found by Galliano et al. (2021).
@@
where $M_{\rm cold}$ is the cold gas mass of the galaxy, $\psi$ is the star-formation rate (SFR), $\beta\psi$ is the rate at which gas is ejected from the cold component into the halo reservoir by SN feedback, $\beta$ is the mass loading factor given by Eq. missing 3, and $R\psi$ is the rate at which gas is returned into the cold component by stellar evolution; the returned fraction, $R$, is determined by the IMF ($R=0.44$ for the Kennicutt 1983 IMF and $R=0.54$ for the tilted, $x=1$, IMF). As galform adopts different IMFs for different SF modes (Section 2.1.2), we calculate the fraction of dust formed ($\tau_{\rm dep}/\tau_{\rm growth}$) for the quiescent and starburst SF modes separately. At high redshift, galaxies have longer dust grain growth timescales because of their lower metallicity, resulting in less dust attenuation than in the original galform dust model.
@@
In Fig. 5 we show predictions for the different assumptions about the dependence of dust mass on metallicity for the EvolFB model (Section 2.1.1), alongside the original dust-extincted (red curve) and the dust-free models (blue curve). The two assumptions, one, that the DTM decreases with decreasing metallicity (the D-Z assumption), and two, that at lower gas metallicity dust takes longer to form (the D-$\tau$ assumption) increase the observed galaxy luminosities and bring the model into better agreement with the data at all redshifts. In both alternative models, the amount of DTM builds up with time and approaches the original model (in which the DTM is independent of redshift and metallicity) at low redshift. At $z=12$, the D-$\tau$ model (green) gives an excellent match to the data, while the D-Z case (orange) only slightly underestimates the abundance at $M_{\rm UV}\sim-20.4$. At $z=14$, the D-$\tau$ model slightly underpredicts the counts of the brightest galaxies.

### v2

@@ 2.1.5 Dust model
Dust attenuates the luminosities of galaxies and is an important consideration when predicting the galaxy luminosity function in different wavebands. galform includes a self-consistent dust model that calculates the absorption of radiation by dust in UV, optical, and near-IR bands, as well as the re-radiation of this energy at IR and sub-mm wavelengths. The model assumes a two-phase dust medium, with dust in molecular clouds embedded in a diffuse dust component. The absorption and re-radiation of energy by the two dust components are calculated separately and, thus, the temperature of the different dust components can vary. Some simplifying assumptions are made: (i) a single dust temperature is assumed for each dust component (cloud or diffuse); (ii) temperature fluctuations for small grains are ignored; (iii) the dust opacity is assumed to be a power law at IR/sub-mm wavelengths, and (iv) polycyclic aromatic hydrocarbon (PAH) features are ignored. We refer readers to Lacey et al. (2016) for a more detailed description of the dust absorption and emission model.
@@ 2.2 Model parameters
Figure 3: galform predictions for the stellar mass function (left) and the star formation rate (SFR) distribution function (right). In each panel, the results from $z=7$ to $z=16$ are shown using different colours, as indicated in the key. The STD model and EvolFB model are shown with dashed and solid curves respectively.
@@
[…]
Figure 5: Upper panel: the redshift evolution of the star formation rate density (SFRD) from $z=19$ down to $z=0$. The results of the EvolFB model and STD model are indicated by solid and dashed curves respectively, with colours representing the SFRD in different star formation (SF) modes (black for the total SFRD, blue for the SFRD in disc/quiescent SF mode, red for the SFRD in starbursts). Lower panel: the integrated stellar mass density at different redshifts, showing both the EvolFB and STD models, along with the separate contributions from disc and starburst SF, as well as the total. Below $z=7$, the SFRD is measured from the P–Millennium simulation and above $z=11$ from the EAGLE-DMO simulation. In between these redshifts, we apply a weight that changes linearly from 1 to 0 (or 0 to 1) between $z=7$ and $z=11$ for P–Millennium (or EAGLE–DMO), allowing the two simulations to join smoothly.
Following the philosophy behind galform, as set out in Cole et al. (2000), Lacey et al. (2016), and other galform papers, we calibrate the model parameters to reproduce a small set of observational constraints, mainly at low redshift, including: (i) the $b_{J}-$ and $K-$band luminosity functions at $z=0$; (ii) the morphological fractions as a function of $r-$band luminosity of galaxies at $z=0$; (iii) the black hole–bulge mass relation at $z=0$; (iv) the number counts and redshift distribution of sub-millimetre galaxies; and (v) UV galaxy luminosity functions at $z=3$ and $z=6$.
Here, we calculate the same models as Cowley et al. (2018). We adopt the same two SN feedback models, STD and EvolFB, described in Section 2.1.2, retaining the same values of all the model parameters. For example, we did not change the parameters of the EvolFB model to reflect a recent change in the inferred redshift of reionization from Planck data, from $z_{\rm re}\gtrsim 8.5$ (Planck Collaboration et al. 2016) to $z_{\rm re}\sim 7-8$ (Planck Collaboration et al. 2020).
The main properties of the local galaxy population and of the submillimetre galaxies used to calibrate the model parameters for both the STD and EvolFB models applied to the P–Millennium simulation are shown in Fig. 2. All other properties of the resulting galaxy population are predictions of the model. In particular, the UV luminosity function of galaxies at the redshifts relevant here, $z=7-16$, was not used to calibrate the model. As we can see from Fig. 2, both the STD model and the EvolFB model show remarkable agreement with the observations at lower redshifts, which is consistent with the results of Lacey et al. (2016) and Cowley et al. (2019) for the STD model and of Hou et al. (2016) for the EvolFB model. The STD model is quite close to the EvolFB model at $z=0$ as the evolving SN feedback converges to the standard case below $z=4$ (Hou et al., 2016).
@@
3.1 Galaxy properties
Before comparing galform predictions with the observational data from JWST, in Fig. 3 we present the predictions of the galform model for (i) the stellar mass function (SMF) and (ii) the star formation rate (SFR) distribution function. Here, we also combine the P–Millennium and EAGLE–DMO simulations using the method described in Section 2.1.1 in order to mitigate the effects of the low mass resolution of P–Millennium. We can see from the left panel of Fig. 3 that the stellar mass function evolves significantly from $z=7$ to $z=16$. The SMF at the lower redshifts shows both a higher number density at fixed stellar mass and extends to larger masses. Across the whole redshift and stellar mass ranges, the EvolFB model predicts a higher galaxy abundance than the STD model. This is a natural consequence of the redshift-dependent supernova feedback model, in which the feedback becomes weaker at higher redshifts ($z>4$), allowing star formation to continue more efficiently.
The SFR distribution function (right panel) also shows significant evolution, with the SFR increasing, on average, towards lower redshifts. The SFR distribution function in the EvolFB model has a similar shape to that in the STD model at $z\lesssim 10$. The highest SFRs at $z\gtrsim 12$ hardly exceed $40\,h^{-1}\,\mathrm{M_{\odot}\,yr^{-1}}$ in either model. The SFR distributions here are roughly consistent with those of Cowley et al. (2018, Fig. 3), but we do see slight differences, especially at high redshift ($z\sim 14-16$). There are more galaxies with low SFR ($<0.1\,h^{-1}\,\mathrm{M_{\odot}\,yr^{-1}}$) in our work. This is because, with the combination of the P–Millennium and EAGLE–DMO simulations, we are able to resolve galaxies with lower masses at high redshifts.
Fig. 4 shows the correlations between stellar mass, halo mass, star formation rate, and UV magnitude from $z=7$ to $16$. We define:
@@ 3.1 Galaxy properties
As can be seen, both the STD and EvolFB models exhibit an approximate power-law relationship between halo mass and stellar mass below $M_{\rm halo,ct}\sim 10^{11}\,\mathrm{M_{\odot}}$. At higher halo masses, the relation flattens and the scatter increases significantly. At a given halo mass, lower redshift galaxies are more likely to have higher stellar masses. The EvolFB model predicts higher stellar masses than the STD model at $M_{\rm halo,ct}\lesssim 10^{11}\,\mathrm{M_{\odot}}$, but similar stellar masses to the STD model at higher halo masses. For the sample used in this figure, $M_{\rm UV}<-15$, nearly all galaxies are centrals, so the figure would be very similar had we plotted $M_{\rm halo}$ rather than $M_{\rm halo,ct}$.
@@
The correlations between $M_{\rm UV}$ and other properties exhibit a strong redshift dependence in both models. Stellar mass, star formation rate, and halo mass all show a positive correlation with the UV luminosity of the galaxy, as expected. At fixed stellar mass and halo mass, high-redshift galaxies appear to be brighter in the UV than low-redshift galaxies. This shows that the “bright” galaxies observed at high redshifts are not “massive” galaxies, contrary to what is often stated in the observational literature (e.g., Harikane et al., 2023; Labbé et al., 2023). Interestingly, in the EvolFB model, high-redshift galaxies show a steeper SFR-$M_{\rm UV}$ relation than low-redshift galaxies (see the bottom left panel of Fig. 4), while in the STD model, this feature is not as obvious.
Fig. 5 shows the redshift evolution of the star formation rate density (SFRD; upper panel) and the integrated stellar mass density (lower panel). The integrated stellar mass density is calculated by integrating the SFRD from $z=19$ to a given redshift and then multiplying by a factor $(1-R)$ to obtain the mass in long-lived stars and stellar remnants. Here, $R$ represents the recycled fraction corresponding to the assumed IMF: $R=0.44$ for the Kennicutt (1983) IMF in disc quiescent SF mode and $R=0.54$ for the $x=1$ IMF in starbursts. The total stellar mass density is the sum of the stellar mass densities formed in quiescent and starburst SF modes.
As we have seen, the P–Millennium simulation does not resolve low-mass galaxies at high redshift and thus underestimates the SFRD at those redshifts. In this figure, we show the results of P–Millennium below $z=7$ and EAGLE–DMO above $z=11$. In between, we apply a weight that linearly varies from 1 to 0 (or 0 to 1) between $z=7$ and $z=11$ for P–Millennium (or EAGLE–DMO), enabling the two simulations to join smoothly. As shown, the SFRDs of both the EvolFB model and the STD model peak at $z\sim 2-4$, with the EvolFB model (solid curve) consistently showing a higher SFRD than the STD model at $z\gtrsim 3$. At $z\lesssim 2.5$, the star formation rate in quiescent mode (blue curves) is significantly higher than that in starbursts (red curves) in both models. However, at higher redshifts ($z\gtrsim 2.5$), the EvolFB model exhibits a significantly higher SFRD in starbursts (red solid) compared to discs (blue solid), while the STD model shows only a slight excess in SFRD in starbursts (red dashed) at $2.5\lesssim z\lesssim 9$. This figure confirms that the SFRD continues to decrease towards higher redshifts starting from $z\sim 2-4$ and does not reach zero, at least not before $z\sim 19$.
3.2 UV luminosity functions
Figure 6: An example of how we combine the P–Millennium and EAGLE–DMO simulations to predict the UV luminosity functions. In each panel, the result from P–Millennium is shown by the blue solid curve, that from EAGLE–DMO by the red solid curve, and the combined result by the black dashed curve.
@@ 3.2 UV luminosity functions
As mentioned in Section 2.1.1, in the following sections we combine the P–Millennium and EAGLE–DMO simulations to estimate the UVLFs. This combination allows us not only to study rare bright galaxies at high redshift (thanks to the large volume of P–Millennium) but also to examine the abundance of faint galaxies (thanks to the high mass resolution of EAGLE–DMO). In Fig. 6, we show an example of the effect of combining the two simulations at different redshifts (for simplicity, only results based on the EvolFB model are shown). P–Millennium and EAGLE–DMO are in good agreement at $z\lesssim 10$, and so the combination does not significantly affect the UVLFs at these redshifts, except for the fact that EAGLE–DMO suffers from incompleteness at the brightest end. At higher redshifts, P–Millennium shows a lower abundance at the faint end than EAGLE–DMO due to its lower mass resolution. The combined results agree well with EAGLE–DMO at the faint end and P–Millennium at the bright end. We note that, due to the very limited sample of EAGLE–DMO galaxies at $z\gtrsim 14$, the UVLFs at these redshifts exhibit noticeable fluctuations.
Fig. 7 presents the dust-extincted (red curves) UVLFs (Section 2.1.5) predicted at $z=7-16$ by galform based on the STD model (dashed curves) and the EvolFB model (solid curves). These are compared with observational data at the corresponding redshifts, including the new JWST observations, as well as earlier estimates using other telescopes such as the Hubble Space Telescope (HST). We use different symbols to denote different observational datasets: (i) UVLFs from spectroscopically confirmed ( $z_{\rm spec}$) high-$z$ candidates from JWST are shown using filled green triangles (Harikane et al., 2024), (ii) UVLFs derived using photometric redshifts $(z_{\rm photo})$ from JWST observations are shown using open green triangles (Adams et al., 2023; Bouwens et al., 2023; Castellano et al., 2023; Donnan et al., 2023; Finkelstein et al., 2023; Harikane et al., 2023; Leung et al., 2023; Morishita & Stiavelli, 2023; Pérez-González et al., 2023; Robertson et al., 2023; Willott et al., 2023), while (iii) UVLFs constructed from earlier, pre-JWST $z_{\rm photo}$ data are indicated by open black triangles (Bouwens et al., 2011, 2015, 2021; Bowler et al., 2014, 2020; Finkelstein et al., 2015; McLure et al., 2013; Morishita et al., 2018; Oesch et al., 2014, 2018; Schenker et al., 2013; Stefanon et al., 2019).
@@
As we can see from Fig. 7, the EvolFB model (solid curve) always predicts a higher abundance of galaxies at the faint end across the whole redshift range ($z=7-16$). This is as expected because the evolving supernova feedback model gives weaker feedback at low masses as well as at higher redshifts (Hou et al., 2016). Below $z\sim 12$, the EvolFB model predicts slightly fewer bright galaxies than the STD model. We note that the discrepancy between the STD and EvolFB models at the bright end at $z\lesssim 12$ is slightly larger than in Cowley et al. (2018). This is because of the difference in the dust attenuation models used in this work and in Cowley et al. (2018). Our main conclusions are unaffected by this difference. We note that the effects of dust extinction on the bright end of the UVLF are predicted to be large even at the high redshifts shown here, but we return to this issue in Section 4.1.
@@
At higher redshifts, however, a discrepancy between the models (with dust extinction; red curves) and the observational data becomes apparent. At $z=12$, the STD model (red dashed curve) underpredicts the abundance of galaxies over the entire $M_{\rm UV}$ range, while the EvolFB model (red solid curve) agrees with the data at the faint end ($M_{\rm UV}\gtrsim-18.5$) but underpredicts the galaxy abundance at the bright end ($M_{\rm UV}\lesssim-19$) by $1-2$ dex. At $z=14$, the predictions of the EvolFB model are within the uncertainty of the data at $M_{\rm UV}>-18$, but are significantly lower at the brighter end ($M_{\rm UV}<-18$) by $2-3$ dex. These results agree with the findings of Kannan et al. (2023) who compared UVLFs from JWST with a suite of simulations based on the IllustrisTNG galaxy formation model. Our results indicate that a revision of the original galform model is required to match the observational data. This is the focus of the next section.
@@ 4 Updating galform
As we have just seen, the predictions made following Cowley et al. (2018), based on two different variations of the galform semi-analytic model within the $\Lambda$CDM framework (the STD and EvolFB models; Section 2), are in good agreement with the latest JWST UVLF data for $z\lesssim 10$. However, the models underpredict the abundance of galaxies at the bright end of the UVLF ($M_{\rm UV}<-18$) at higher redshifts ($z\sim 12-14$) when dust extinction is included. This discrepancy could be due to several factors. Here, we explore the sensitivity of the predicted UVLFs to the details of specific physical components of the model – most notably, the dust model and the IMF (described in Section 2). We investigate whether reasonable adjustments of the model parameters make it possible to bring the galform predictions into closer agreement with the observed UVLFs at $z\gtrsim 12$.
@@ 4.1 The effect of dust attenuation
One way to resolve the difference between our model predictions and the observed UVLFs at high redshifts is to alter the importance of dust attenuation, since dust in galaxies absorbs light and thus reduces the number counts of bright galaxies. As a limiting case, we begin by turning off the dust attenuation in galform altogether (for both the STD and EvolFB models) and calculate dust-free UVLFs at $z=7-16$; these predictions are also shown in Fig. 7 (blue curves).
@@
There is some observational evidence that dust extinction may be less important at very high redshift than in the local Universe (e.g. Austin et al. 2024; Cullen et al. 2024; Morales et al. 2024; Topping et al. 2024). Several studies (e.g. Vladilo et al., 2011; De Cia et al., 2013, 2016; Wiseman et al., 2017) have reported a metallicity-dependent dust-to-metal ratio (DTM), which is about 50% lower at metallicity $Z=0.1Z_{\odot}$ than at solar metallicity, $Z_{\odot}$ (see Maiolino & Mannucci 2019 for a review). Since metallicities are lower at high redshift, this would be expected to lead to a lower DTM. Contrary to this, Wiseman et al. (2017) reported that they did not find any obvious redshift-dependence of the DTM; however, this conclusion is based on small galaxy samples up to a redshift of only $z\sim 5$. In galform, a constant DTM is assumed with a value of 0.334 (Lacey et al., 2016) in order to match the dust-to-gas ratio in the solar neighbourhood, $6.7\times 10^{-3}$ for $Z_{\odot}=0.02$ (Silva et al., 1998). If the reports of a lower DTM ratio at lower metallicity are correct, galform would overestimate the dust content at low metallicity and, consequently, that of galaxies at high redshifts.
Another, more radical, possibility is that dust may not have had time to form, in some of the high redshift galaxies (e.g. Markov et al. 2024). The details of dust formation and destruction mechanisms are very uncertain even at low redshift but, from analytical and observational considerations, Mattsson & Andersen (2012) concluded that dust growth in the interstellar medium is the most important dust formation mechanism. They proposed a dust grain growth timescale, $\tau_{\rm growth}$, which is found to increase with decreasing metallicity and can be as long as $\gtrsim 1\,\mathrm{Gyr}$ for the lowest metallicity galaxies (Galliano et al., 2021). Since $z=14$ corresponds to only $\sim 300\,\mathrm{Myrs}$ after the Big Bang, galaxies at this redshift may not have had enough time to make a significant amount of dust. Schneider & Maiolino (2024) proposed a parametrization of $\tau_{\rm growth}$ which depends on the distribution of grain sizes, and on the density, temperature, and metallicity of the interstellar medium. This model also shows a negative correlation with metallicity, similar to that found by Galliano et al. (2021).
@@
where $M_{\rm cold}$ is the cold gas mass of the galaxy, $\psi$ is the star formation rate (SFR), $\beta\psi$ is the rate at which gas is ejected from the cold component into the halo reservoir by SN feedback, $\beta$ is the mass loading factor given by Eq. missing 3, and $R\psi$ is the rate at which gas is returned into the cold component by stellar evolution; the returned fraction, $R$, is determined by the IMF ($R=0.44$ for the Kennicutt 1983 IMF and $R=0.54$ for the tilted, $x=1$, IMF). As galform adopts different IMFs for different SF modes (Section 2.1.3), we calculate the fraction of dust formed ($\tau_{\rm dep}/\tau_{\rm growth}$) for the quiescent and starburst SF modes separately. At high redshift, galaxies have longer dust grain growth timescales because of their lower metallicity, resulting in less attenuation than in the original galform dust model.
@@
In Fig. 8 we show predictions for the different assumptions about the dependence of dust mass on metallicity for the EvolFB model (Section 2.1.2), alongside the original dust-extincted (red curve) and the dust-free models (blue curve). The two assumptions, one, that the DTM decreases with decreasing metallicity (the D-Z assumption), and two, that at lower gas metallicity dust takes longer to form (the D-$\tau$ assumption) increase the observed galaxy luminosities and bring the model into better agreement with the data at all redshifts. In both alternatives, the amount of DTM builds up with time and approaches the original model (in which the DTM is independent of redshift and metallicity) at low redshift. At $z=12$, the D-$\tau$ model (green) gives an excellent match to the data, while the D-Z case (orange) only slightly underestimates the abundance at $M_{\rm UV}\sim-20.4$. At $z=14$, the D-$\tau$ model slightly underpredicts the counts of the brightest galaxies.